OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

OpenStudy (ajprincess):

when u subtract the 2nd eqn from 1st u get 2x=2 x=1 Applyin t in the 1st eqn -2+y=1 y=3 (1,3)

OpenStudy (anonymous):

Rearrange -2x + y = 1 y = 2x + 1 sub y = 2x + 1 into second equation -4x + 2x + 1 = -1 -2x = -2 x = 1 Sub x = 1 back into either equation one or two lets sub it back into equation one. -2(1) + y = 1 -2 + y = 1 so y = 3 Now u have ur x = 1 and y = 3 so it is (1,3)